# Towards a generic absoluteness theorem for Chang models

### v2

@@ 5.4. Summary
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground, $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$” and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause 3 of Definition 4.7 applied to $(X,W)$. Suppose $k\subseteq\operatorname{Col}(\omega,{|\Gamma^{\infty}|})$ is $V$-generic, $\iota\leq\omega_{1}^{V}$ is a limit ordinal, $k^{\prime}\in V[k]$ is a surjection $k^{\prime}:\iota\rightarrow\Gamma^{\infty}$ with the property that for all $\alpha<\iota$, $k^{\prime}\restriction\alpha\in V$ and
@@
be the underlying sequence of ${\sf{DM}}$, $M_{\iota}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\iota}:M_{\iota}\rightarrow U_{\chi}$ be the canonical realizable embedding. Then in $V[k]$, there is an $M_{\iota}$-generic $h\subseteq\operatorname{Col}(\omega,{<}\omega_{1}^{V})$ such that
@@
Notice that we have that $\kappa_{\iota}=\omega_{1}^{V}$. The existence of $h$ follows from [6, Lemma 3.1.5]. Because $\sigma_{\iota}:M_{\iota}\rightarrow U_{\chi}$ and there are Woodin cardinals in $U_{\chi}$ (see Clause 3 of Definition 4.7), $M_{\iota}[h]\models``(Hom^{*},\mathbb{R}^{*})^{\#}$ exists"""" and $((Hom^{*},\mathbb{R}^{*})^{\#})^{M_{\iota}[h]}=(Hom^{*},\mathbb{R}^{*})^{\#}$. It is then enough to prove that if $h\subseteq\operatorname{Col}(\omega,{<}\omega_{1}^{V})$ is $M_{\iota}$-generic such that $h\in V[k]$ and $\mathbb{R}^{M_{\iota}[h]}=\mathbb{R}^{V}$ then
@@
First fix $A\in\Gamma^{\infty}$. We want to show that $A\in(Hom^{*})^{M_{\iota}[h]}$. Let $\alpha<\iota$ be such that ${\sf a}_{\alpha}$ b-catches $A$ relative to $(U,\chi,\eta,\lambda,W^{\prime})$. Let $\beta<\kappa_{\alpha}$ and $f\subseteq\operatorname{Col}(\omega,\beta)$ be such that $f\in V$, $f$ is $M_{\alpha}$-generic and there is $\bar{\mu}\in M_{\alpha}[f]$ such that $\bar{\nu}=_{def}\sigma_{\alpha}\,"\bar{\mu}\subseteq W^{\prime}$ and $S_{\bar{\nu}}=A$. We can then find some $\gamma\in[\alpha,\iota)$ such that $f\in M_{\gamma}[h_{\gamma}]$. Because we have that $\sigma_{\alpha,\gamma}$ lifts to $\sigma_{\alpha,\gamma}^{f}:M_{\alpha}[f]\rightarrow M_{\gamma}[f]$ and because $M_{\gamma}[f]\subseteq M[h_{\gamma}]$, we have that letting $\bar{\tau}=\sigma_{\alpha,\gamma}^{+}\,"\bar{\mu}$,
@@
(2) $A=\bigcap_{\xi\in(\gamma,\iota)}A_{\xi}$ and $A_{\xi}=(S_{\bar{\tau}_{\xi}})^{M_{\xi}[h_{\xi}]}$ (see Lemma 3.2).
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground, $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$” and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause 3 of Definition 4.7 applied to $(X,W)$. Suppose $k\subseteq\operatorname{Col}(\omega_{1},{|\Gamma^{\infty}|})$ is $V$-generic and
@@
be the underlying sequence of ${\sf{DM}}$, $M_{\omega_{1}}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\omega_{1}}:M_{\omega_{1}}\rightarrow U_{\chi}$ be the canonical realizable embedding. Then in $V[k]$, there is an $M_{\omega_{1}}$-generic $h\subseteq\operatorname{Col}(\omega,{<}\omega_{1}^{V})$ such that
@@
This remark constructs the generic $h$ used in Theorem 6.6. Notice that because we are working in a universe that has class of Woodin cardinals, countably closed posets do not add new universally Baire sets as they do not add new homogeneity systems (see Theorem 2.3). Let $\vec{x}=(x_{i}\mid i<\alpha)\in V[k]$ be an enumeration of $\mathbb{R}(=\mathbb{R}^{V})$ and let $f:\omega_{1}\rightarrow\omega_{1}$ be any increasing continuous function in $V[k]$ such that for every $\beta<\omega_{1}$ there is an $M_{f(\beta)}$-generic $g\subseteq\operatorname{Col}(\omega,{<}\kappa_{f(\beta)})$ such that $x_{\beta}\in M_{f(\beta)}[g]$. Let $\mathbb{P}$ consist of functions $p$ such that
@@
    for each $\beta\in dom(p)$, $p(\beta)$ is an $M_{f(\beta)}$-generic for $\operatorname{Col}(\omega,{<}\kappa_{f(\beta)})$ such that $\vec{x}\restriction\beta\subseteq M_{f(\beta)}[p(\beta)]$,
@@
The order of $\mathbb{P}$ is by extension. It follows from the proof of Claim 3 that appears in the proof of [6, Lemma 3.1.5] that each $p\in\mathbb{P}$ has a proper extension in $\mathbb{P}$. Notice now that $\mathbb{P}$ is countably closed [^42] This is because if $(p_{i}:i<\omega)$ is a decreasing sequence of conditions then letting $\beta=\sup\{\gamma:\exists i\in\omega(\gamma\in dom(p_{i}))\}$ and $g=\cup_{i<\omega}p_{i}(\beta_{i})$ where $\beta_{i}$ is the largest element of $dom(p_{i})$, $g\subseteq\operatorname{Col}(\omega,{<}\kappa_{\beta})$ is $M_{\beta}$-generic.. Let then $\xi$ be large and $\nu:H\rightarrow V_{\xi}[k]$ be an elementary embedding such that in $V[k]$, $H$ is transitive, ${|H|}=\aleph_{1}$, $H^{\omega}\subseteq H$ and $\{\Gamma^{\infty},\mathbb{P},D\}\subseteq\operatorname{rng}(\nu)$. We now have $h^{\prime}\in V[k]$ which is $H$-generic for $\mathbb{P}$. Let then $h=\bigcup_{p\in h^{\prime},\beta\in dom(p)}p(\beta)$. It follows that $h$ is $M_{\omega_{1}}$-generic for $\operatorname{Col}(\omega,{<}\omega_{1})$.
@@ 7. $\mathsf{Sealing}$ from the derived model representation
In this section, we use the derived model representation in Theorem 6.5 to prove $\mathsf{Sealing}$ similarly as in [22, Section 6].
@@
We now work towards the set up of Theorem 4.10, and so we work in $V$. Let $\lambda$ be an inaccessible cardinal such that $g*h*h^{\prime}$ is a $<\lambda$-generic over $V$, $\chi>\lambda$ be an inaccessible cardinal, $W\in V_{\chi}$ a $(\chi,\lambda^{+},\lambda)$-saturated set of measures with the property that there are infintely many Woodin cardinals in the interval, $X\prec V_{\chi}$ be such that ${|X|}={|V_{\kappa+1}|}$ and $V_{\kappa+1}\cup\{\lambda,W\}\subseteq X$ and let $j:V\rightarrow N$ be a $\chi$-supercompactness embedding. We now have that the UB-Capturing Principle holds in $N$ at $(j(\chi),j(\kappa),\lambda)$ as witnessed by $(j[W],j[X\cap V_{\chi}])$. In what follows we apply the lemmas and theorems of the previous sections to $(V,N[g*h])$ and $(V,N[g*h*h^{\prime}])$ with $V$ as the role of $U$. In particular, we apply Theorem 6.5 with $U$ as $V$, $\chi$ as $j(\chi)$, $\eta$ as $j(\kappa)$, $\lambda$ as $\lambda$, $X$ as $j[X\cap V_{\chi}]$, $W$ as $W$ and $W^{\prime}$ as $j(W)$.
@@
Let $k\subseteq\operatorname{Col}(\omega,{|\Gamma^{\infty}_{g*h}|})$ be an $N[g*h]$-generic and $k^{\prime}\subseteq\operatorname{Col}(\omega,{|\Gamma^{\infty}_{g*h*h^{\prime}}|})$ be an $N[g*h*h^{\prime}]$-generic such that $k(0)=A$ and $k^{\prime}(0)=A_{h}$. Applying Theorem 6.5 in $N[g*h]$, we obtain a ${\sf{DM}}$-sequence
@@
Let $l_{1}=l\cap\operatorname{Col}(\omega,{<}\kappa_{1})$ and $l_{1}^{\prime}=l\cap\operatorname{Col}(\omega,{<}\kappa_{1}^{\prime})$. We then have that
@@ 8. $\mathsf{Sealing}$ for derived models
    For $A\in(\operatorname{Hom}^{*})^{V[g*H]}$, say $A=B^{*}$ for some $B\in\operatorname{Hom}_{{<}\eta}^{V[g*H\upharpoonright\gamma]}$, then $k(A)$ is the canonical interpretation of $B$ in $V[g*H]$.
@@
Note that $(L(\operatorname{Hom}^{*},\mathbb{R}^{*}))^{V[g*H]}$ is equal to the direct limit of $L(\Gamma^{\infty}_{g*H_{i}},\mathbb{R}_{g*H_{i}})$ under the elementary embeddings $j_{i,k}$ for $i<k<\operatorname{cf}(\eta)$. Indeed, if $A\in(L(\operatorname{Hom}^{*},\mathbb{R}^{*}))^{V[g*H]}\cap({\wp}(\mathbb{R}^{*}))^{V[g*H]}$ then, for some $i$ and some $B\in\Gamma^{\infty}_{g*H_{i}}$, $A=B^{*}$. Let
@@
such that $i<\operatorname{cf}(\eta)$, $j_{i}=j_{\infty}\circ j_{i,\infty}.$
We are left with showing the following claim.
@@ 9. Generically Correct Sealing
    Whenever $g$ is $<\lambda$-generic and $\bar{\mu}\subseteq W$ is in $V[g]$, $S_{\bar{\mu}_{g}}\in\Gamma^{\infty}_{g}$.
@@
    such that for every $\bar{\mu}\subseteq\pi_{X}^{-1}(W)$ with $\bar{\mu}\in N_{X}[G]$,
@@
Let $\lambda<\chi$ be inaccessible cardinals in $V[g]$ and $W_{0}\in V_{\chi}[g]$ be a $(\chi,\lambda^{+},\lambda)$-saturated and infinitely tamed set of measures. We can then find $W\in V$ such that $W_{0}\subseteq W$ and in $V$ (and hence in $V[g]$), $W$ is a $(\chi,\lambda^{+},\lambda)$-saturated and infinitely tamed set of measures. It is enough to verify the claim for $(\lambda,\chi,W)$.
Clause (1) follows from Lemma 2.11 and is also proven in [6] (e.g. see [6, Lemma 3.4.11]). We verify Clause (2). Let $X\prec V_{\chi}$ be a countable substructure. Let $\pi_{X}\colon N_{X}\rightarrow V_{\chi}[g]$ be the Mostowski collapse of $X$. In what follows we will use subscript $X$ to denote the $\pi_{X}$-preimages of various objects, and we will use this convention with other substructures as well. Let $G$ be a ${<}\lambda_{X}$-generic over $N_{X}$. Then the elementary embedding
@@
(a) For a club of $X\prec V_{\chi}[g]$, if $G$ is a $<\lambda_{X}$-generic, $\bar{\mu}\subseteq W_{X}$ with $\bar{\mu}\in N_{X}[G]$, $\phi$ is a formula and $c_{0},...,c_{n}$ are constants for indiscernibles, then
@@
(b) For all $X\prec V_{\theta^{\prime}}[g]$ such that $\{Y,\chi,\lambda,\theta,j\restriction V_{\theta}\}\in X$, if $G$ is a $<\lambda_{X}$-generic over $N_{X}[G]$, $\bar{\mu}\subseteq W_{X}$ with $\bar{\mu}\in N_{X}[G]$, $\phi$ is a formula and $c_{0},...,c_{n}$ are constants for indiscernibles, then
(a) follows from (b) because the set of $X^{\prime}\prec V_{\chi}[g]$ such that for some $X$ as in (b), $X^{\prime}=X\cap V_{\chi}[g]$ is a club. Fix then an $X$ as in (b), and and also fix $G\in V[g]$ that is a $<\lambda_{X}$-generic over $N_{X}$, $\bar{\mu}\subseteq W_{X}$ with $\bar{\mu}\in N_{X}[G]$, a formula $\phi$ and $c_{0},...,c_{n}$ constants for indiscernibles. We set $P=j(V_{\theta})$. Following the proof of Theorem 1.9, we can find, in $N_{X}[G]$, a $(P_{X},j_{X}(\chi),\lambda_{X},\{j_{X}(\kappa),j_{X}(W_{X})\})$-realizable rock [^43] See Definition 5.5. We are abusing terminology here, as $P_{X}$ is not a class, but clearly this is irrelevant. ${\sf a}$ and $k\subseteq\operatorname{Col}(\omega,{<}\kappa^{{\sf a}})$ such that for some $\bar{\nu}\in M^{\sf a}[k]$,
(1.1) in $N_{X}[G]$, ${\sf a}$ b-catches $(S_{\bar{\mu}})^{N_{X}[G]}$ relative to $(P_{X},j_{X}(\chi),j_{X}(\kappa),\lambda_{X},j_{X}(W_{X}))$ (see Definition 5.14),
@@
(2.1) ${\sf{b}}$ b-catches $S_{\pi_{X}\circ\sigma^{\sf a}\,"\bar{\nu}}$ relative to $(P,j(\chi),j(\kappa),\lambda,j(W_{X}))$,
@@
(2.3) $\pi_{X}\circ\sigma^{\sf a}\,"\bar{\nu}\subseteq j(W)$ and
@@
Assume there is a proper class of inaccessible limit of Woodin cardinals. Then $\mathsf{Generically}\ \allowbreak\mathsf{Correct}\ \allowbreak\mathsf{Sealing}$ implies $\mathsf{Sealing}$.
@@
such that for every $A\in\Gamma^{\infty}_{h_{1}}$, $j(A)=A_{h_{2}}$. The derived model theorem (see Theorem 2.2) implies that assuming $\mathsf{Sealing}$ and a class of inaccessible limits of Woodin cardinals, Clause (2) implies Clause (1).

### v3

@@ 5.4. Summary
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground, $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$” and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause (3) of Definition 4.7 applied to $(X,W)$. Suppose $k\subseteq\operatorname{Col}(\omega,{|\Gamma^{\infty}|})$ is $V$-generic, $\iota<\omega_{1}^{V}$ is a limit ordinal, $k^{\prime}\in V[k]$ is a surjection $k^{\prime}:\iota\rightarrow\Gamma^{\infty}$ with the property that for all $\alpha<\iota$, $k^{\prime}\restriction\alpha\in V$ and
@@
be the underlying sequence of $D$, $M_{\iota}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\iota}:M_{\iota}\rightarrow U_{\chi}$ be the canonical realizability embedding. Then, in $V[k]$, there is an $M_{\iota}$-generic $h\subseteq\operatorname{Col}(\omega,{<}\omega_{1}^{V})$ such that
@@
Notice that we have that $\kappa_{\iota}=\omega_{1}^{V}$. The existence of $h$ follows from [7, Lemma 3.1.5]. Because $\sigma_{\iota}:M_{\iota}\rightarrow U_{\chi}$ and there are Woodin cardinals in $U_{\chi}$ (see Clause (3) of Definition 4.7), $M_{\iota}[h]\models``(Hom^{*},\mathbb{R}^{*})^{\#}$ exists"""" and $((Hom^{*},\mathbb{R}^{*})^{\#})^{M_{\iota}[h]}=(Hom^{*},\mathbb{R}^{*})^{\#}$. It is then enough to prove that if $h\subseteq\operatorname{Col}(\omega,{<}\omega_{1}^{V})$ is $M_{\iota}$-generic such that $h\in V[k]$ and $\mathbb{R}^{M_{\iota}[h]}=\mathbb{R}^{V}$ then
@@
First fix $A\in\Gamma^{\infty}$. We want to show that $A\in(Hom^{*})^{M_{\iota}[h]}$. Let $\alpha<\iota$ be such that ${\sf a}_{\alpha}$ b-catches $A$ relative to $(U,\chi,\lambda,\{\eta,W^{\prime}\})$. Let $\beta<\kappa_{\alpha}$ and $f\subseteq\operatorname{Col}(\omega,\beta)$ be such that $f\in V$, $f$ is $M_{\alpha}$-generic and there is $\bar{\mu}\in M_{\alpha}[f]$ such that $\bar{\nu}=_{def}\sigma_{\alpha}\,"\bar{\mu}\subseteq W^{\prime}$ and $S_{\bar{\nu}}=A$. We can then find some $\gamma\in[\alpha,\iota)$ such that $f\in M_{\gamma}[h_{\gamma}]$. Because we have that $\sigma_{\alpha,\gamma}$ lifts to $\sigma_{\alpha,\gamma}^{f}:M_{\alpha}[f]\rightarrow M_{\gamma}[f]$ and because $M_{\gamma}[f]\subseteq M[h_{\gamma}]$, we have that letting $\bar{\tau}=\sigma_{\alpha,\gamma}^{+}\,"\bar{\mu}$,
@@
(2) $A=\bigcup_{\xi\in(\gamma,\iota)}A_{\xi}$ and $A_{\xi}=(S_{\bar{\tau}_{\xi}})^{M_{\xi}[h_{\xi}]}$ (see Lemma 3.2).
@@
Suppose $\chi>\eta>\lambda$ are inaccessible cardinals, $U$ is a $\lambda$-ground, $U\models``$the UB-Capturing Principle holds at $(\chi,\eta,\lambda)$ as witnessed by $(X,W)$” and $W^{\prime}\in X$ is such that $W\subseteq W^{\prime}$ and $W^{\prime}$ witnesses Clause (3) of Definition 4.7 applied to $(X,W)$. Suppose $k\subseteq\operatorname{Col}(\omega_{1},{|\Gamma^{\infty}|})$ is $V$-generic and
@@
be the underlying sequence of $D$, $M_{\omega_{1}}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\omega_{1}}:M_{\omega_{1}}\rightarrow U_{\chi}$ be the canonical realizability embedding. Then, in $V[k]$, there is an $M_{\omega_{1}}$-generic $h\subseteq\operatorname{Col}(\omega,{<}\omega_{1}^{V})$ such that
@@
This remark constructs the generic $h$ used in Theorem 6.6. Notice that because we are working in a universe that has a class of Woodin cardinals, countably closed posets do not add new universally Baire sets as they do not add new homogeneity systems (see Theorem 2.3). Let $\vec{x}=(x_{i}\mid i<\alpha)\in V[k]$ be an enumeration of $\mathbb{R}(=\mathbb{R}^{V})$ and let $f:\omega_{1}\rightarrow\omega_{1}$ be any increasing continuous function in $V[k]$ such that for every $\beta<\omega_{1}$ there is an $M_{f(\beta)}$-generic $g\subseteq\operatorname{Col}(\omega,{<}\kappa_{f(\beta)})$ such that $x_{\beta}\in M_{f(\beta)}[g]$. Let $\mathbb{P}$ consist of functions $p$ such that
@@
    for each $\beta\in dom(p)$, $p(\beta)$ is $M_{f(\beta)}$-generic for $\operatorname{Col}(\omega,{<}\kappa_{f(\beta)})$ such that $\vec{x}\restriction\beta\subseteq M_{f(\beta)}[p(\beta)]$,
@@
The order of $\mathbb{P}$ is by extension. It follows from the proof of Claim 3 that appears in the proof of [7, Lemma 3.1.5] that each $p\in\mathbb{P}$ has a proper extension in $\mathbb{P}$. Notice now that $\mathbb{P}$ is countably closed [^45] This is because if $(p_{i}:i<\omega)$ is a decreasing sequence of conditions then letting $\beta=\sup\{\gamma:\exists i\in\omega(\gamma\in dom(p_{i}))\}$ and $g=\bigcup_{i<\omega}p_{i}(\beta_{i})$ where $\beta_{i}$ is the largest element of $dom(p_{i})$, $g\subseteq\operatorname{Col}(\omega,{<}\kappa_{\beta})$ is $M_{\beta}$-generic.. Let then $\xi$ be large and $\nu:H\rightarrow V_{\xi}[k]$ be an elementary embedding such that in $V[k]$, $H$ is transitive, ${|H|}=\aleph_{1}$, $H^{\omega}\subseteq H$ and $\{\Gamma^{\infty},\mathbb{P},D\}\subseteq\operatorname{rng}(\nu)$. We now have $h^{\prime}\in V[k]$ which is $H$-generic for $\mathbb{P}$. Let then $h=\bigcup_{p\in h^{\prime},\beta\in dom(p)}p(\beta)$. It follows that $h$ is $M_{\omega_{1}}$-generic for $\operatorname{Col}(\omega,{<}\omega_{1})$.
@@ 7. $\mathsf{Sealing}$ from the derived model representation
In this section, we use the derived model representation in Theorem 6.5 to prove $\mathsf{Sealing}$ similarly as in [21, Section 6].
@@
We now work towards the set up of Theorem 4.10, and so we work in $V$. Let $\lambda$ be an inaccessible cardinal such that $g*h*h^{\prime}$ is ${<}\lambda$-generic over $V$ and let $\chi>\lambda$ be an inaccessible cardinal. Let $W\in V_{\chi}$ be a $(\chi,\lambda^{+},\lambda)$-saturated set of measures with the property that there are infinitely many Woodin cardinals in the interval $(\lambda,{\sf{comp}}(W))$, let $X\prec V_{\chi}$ be such that ${|X|}={|V_{\kappa+1}|}$ and $V_{\kappa+1}\cup\{\lambda,W\}\subseteq X$ and let $j:V\rightarrow N$ be a $\chi$-supercompactness embedding. We now have that the UB-Capturing Principle holds in $N$ at $(j(\chi),j(\kappa),\lambda)$ as witnessed by $(j[W],j[X\cap V_{\chi}])$. In what follows we apply the lemmas and theorems of the previous sections to $(V,N[g*h])$ and $(V,N[g*h*h^{\prime}])$ with $V$ in the role of $U$. In particular, we apply Theorem 6.5 with $U$ as $V$, $\chi$ as $j(\chi)$, $\eta$ as $j(\kappa)$, $\lambda$ as $\lambda$, $X$ as $j[X\cap V_{\chi}]$, $W$ as $W$ and $W^{\prime}$ as $j(W)$.
@@
Let $k\subseteq\operatorname{Col}(\omega,{|\Gamma^{\infty}_{g*h}|})$ be $N[g*h]$-generic and $k^{\prime}\subseteq\operatorname{Col}(\omega,{|\Gamma^{\infty}_{g*h*h^{\prime}}|})$ be $N[g*h*h^{\prime}]$-generic such that $k(0)=A$ and $k^{\prime}(0)=A_{h}$. Applying Theorem 6.5 in $N[g*h]$, we obtain a ${\sf{DM}}$-sequence
@@
Recall that $\kappa_{1}=\kappa^{{\sf a}_{1}}$ and $\kappa_{1}^{\prime}=\kappa^{{\sf a}_{1}^{\prime}}$. Let $l_{1}=l\cap\operatorname{Col}(\omega,{<}\kappa_{1})$ and $l_{1}^{\prime}=l\cap\operatorname{Col}(\omega,{<}\kappa_{1}^{\prime})$. We then have that
@@ 8. $\mathsf{Sealing}$ for derived models
    For $A\in(\operatorname{Hom}^{*})^{V[g*H]}$, say $A=B^{*}$ for some $B\in\operatorname{Hom}_{{<}\eta}^{V[g*H\upharpoonright\gamma]}$ for some $\gamma<\eta$, then $k(A)$ is the canonical interpretation of $B$ in $V[g*H]$.
@@
Note that
Claim 1.
$(L(\operatorname{Hom}^{*},\mathbb{R}^{*}))^{V(\mathbb{R}^{*})}$ is equal to the direct limit of $L(\Gamma^{\infty}_{g*H_{i}},\mathbb{R}_{g*H_{i}})$ under the elementary embeddings $j_{i,k}$ for $i<k<\operatorname{cf}(\eta)$.
Proof.
Let
@@
such that $i<\operatorname{cf}(\eta)$, $j_{i}=j_{\infty}\circ j_{i,\infty}.$ Moreover, if $X=j_{\infty}(\Gamma^{\infty}_{g*H})$ and $Y=j_{\infty}^{-1}(\mathbb{R}_{g*H})$ then $M=L(X,Y)$.
It is enough to show that $X=Hom^{*}$ and $Y=\mathbb{R}^{*}$. Suppose first that $x\in Y$. We can then find $i<\operatorname{cf}(\eta)$ such that $x\in\operatorname{rng}(j_{i,\infty})$. Then it follows that $x\in\mathbb{R}_{g*H_{i}}$, and hence $x\in\mathbb{R}^{*}$. Similarly, if $x\in\mathbb{R}^{*}$ then for some $i<\operatorname{cf}(\eta)$, $x\in\mathbb{R}_{g*H_{i}}$ and so, $x\in Y$ as $j_{i,\infty}(x)=x$.
Suppose now that $A\in Hom^{*}$. We want to show that $A\in X$. Fix $i<\operatorname{cf}(\eta)$ and some $B\in\Gamma^{\infty}_{g*H_{i}}$ so that $A=B^{*}$. We have for every $k\in[i,\operatorname{cf}(\eta))$, $j_{i,k}(B)=B_{g*H_{k}}$. As $A=\bigcup_{k\in[i,\operatorname{cf}(\eta))}B_{g*H_{k}}$ and $j_{i,\infty}(B)=\bigcup_{k\in[i,\operatorname{cf}(\eta))}j_{i,k}(B)$, we get $A=j_{i,\infty}(B)$. Hence, $A\in X$ as $A=j_{\infty}^{-1}(j_{i}(B))$.
Suppose next that $A\in X$. We have $i<\operatorname{cf}(\eta)$ and $B\in\Gamma^{\infty}_{g*H_{i}}$ such that $A=j_{i,\infty}(B)$. Arguing as above, we get that $A=B^{*}$. ∎
We are left with the following claim (which was implicitly proven above).
@@ 9. Generically Correct Sealing
    Whenever $g$ is ${<}\lambda$-generic and $\bar{\mu}\subseteq W$ is a homogeneity system in $V[g]$, $S_{\bar{\mu}_{g}}\in\Gamma^{\infty}_{g}$.
@@
    such that for every homogeneity system $\bar{\mu}\subseteq\pi_{X}^{-1}(W)$ with $\bar{\mu}\in N_{X}[G]$,
@@
Let $\lambda<\chi$ be inaccessible cardinals in $V[g]$ and $W_{0}\in V_{\chi}[g]$ be a $(\chi,\lambda^{+},\lambda)$-saturated and infinitely tamed set of measures. We can then find $W\in V$ such that $W_{0}\subseteq W$ and in $V$ (and hence in $V[g]$), $W$ is a $(\chi,\lambda^{+},\lambda)$-saturated and infinitely tamed set of measures. It is enough to verify Clauses (1) and (2) in the definition of $\mathsf{Generically}\ \allowbreak\mathsf{Correct}\ \allowbreak\mathsf{Sealing}$ for $(\lambda,\chi,W)$.
Clause (1) follows from Lemma 2.11 and is also proven in [7] (e.g. see [7, Lemma 3.4.11]). We verify Clause (2). Let $X\prec V_{\chi}$ be a countable substructure. Let $\pi_{X}\colon N_{X}\rightarrow V_{\chi}[g]$ be the Mostowski collapse of $X$. In what follows we will use subscript $X$ to denote the $\pi_{X}$-preimages of various objects, and we will use this convention with other substructures as well. Let $G$ be ${<}\lambda_{X}$-generic over $N_{X}$. Then the elementary embedding
@@
(a) For a club of $X\prec V_{\chi}[g]$, if $G$ is ${<}\lambda_{X}$-generic, $\bar{\mu}\subseteq W_{X}$ is a homogeneity system with $\bar{\mu}\in N_{X}[G]$, $\phi$ is a formula and $c_{0},...,c_{n}$ are constants for indiscernibles, then
$(\phi,S_{\bar{\mu}}^{N_{X}[G]},c_{0},...,c_{n})\in((\Gamma^{\infty}_{G},\mathbb{R}_{G})^{\#})^{N_{X}[G]}\iff(\phi,S_{\pi_{X}\,"\bar{\mu}},c_{0},...,c_{n})\in(\Gamma^{\infty},\mathbb{R})^{\#}.$
(b) For all $X\prec V_{\theta^{\prime}}[g]$ such that $\{Y,\chi,\lambda,\theta,j\restriction V_{\theta}\}\in X$, if $G$ is ${<}\lambda_{X}$-generic over $N_{X}[G]$, $\bar{\mu}\subseteq W_{X}$ is a homogeneity system with $\bar{\mu}\in N_{X}[G]$, $\phi$ is a formula and $c_{0},...,c_{n}$ are constants for indiscernibles, then
@@
We first show that (a) follows from (b). Assume (b). Let $T\subseteq\omega\times[V_{\chi}[g]]^{<\omega}$ be the theory of $V_{\theta^{\prime}}$ with parameters from $\{Y,\chi,\lambda,\theta,j\restriction V_{\theta}\}$. More precisely, $T$ consists of pairs $(\phi,\vec{a})$ where $\phi$ is a formula and $\vec{a}$ is a finite sequence of members of $V_{\chi}[g]$ such that
$V_{\theta^{\prime}}[g]\models\phi[\vec{a},Y,\chi,\lambda,\theta,j\restriction V_{\theta}].$
Suppose now that $X^{\prime}\prec(V_{\chi}[g],T,\in)$. We claim that (a) holds for any such $X^{\prime}$. Let $X$ be the Skolem hull of $X^{\prime}\cup\{Y,\chi,\lambda,\theta,j\restriction V_{\theta}\}$ computed in $V_{\theta^{\prime}}[g]$. We claim that $X\cap V_{\chi}[g]=X^{\prime}$. Suppose that $\vec{a}\in[X^{\prime}]^{<\omega}$ and $b\in V_{\chi}[g]$ is such that for some formula $\phi$, $b$ is the unique set such that
$V_{\theta^{\prime}}[g]\models\phi[b,\vec{a},Y,\chi,\lambda,\theta,j\restriction V_{\theta}].$
It then follows that $b$ is the unique member of $V_{\chi}[g]$ such that $(\phi,b^{\frown}\vec{a})\in T$. Since $\vec{a}\in X^{\prime}$, we have that $b\in X^{\prime}$. We can now apply (b) to $X$ and get (a) for $X^{\prime}$.
Fix then an $X$ as in (b), and also fix $G\in V[g]$ that is ${<}\lambda_{X}$-generic over $N_{X}$, a homogeneity system $\bar{\mu}\subseteq W_{X}$ with $\bar{\mu}\in N_{X}[G]$, a formula $\phi$ and $c_{0},...,c_{n}$ constants for indiscernibles. We set $P=j(V_{\theta})$. Following the proof of Theorem 1.9, we can find, in $N_{X}[G]$, a $(P_{X},j_{X}(\chi),\lambda_{X},\{j_{X}(\kappa),j_{X}(W_{X})\})$-realizable rock [^46] See Definition 5.5. We are abusing terminology here as $P_{X}$ is not a class, but clearly this is irrelevant. ${\sf a}$ and $k\subseteq\operatorname{Col}(\omega,{<}\kappa^{{\sf a}})$ such that for some $\bar{\nu}\in M^{\sf a}[k]$,
(1.1) in $N_{X}[G]$, ${\sf a}$ b-catches $(S_{\bar{\mu}})^{N_{X}[G]}$ relative to $(P_{X},j_{X}(\chi),\lambda_{X},\{j_{X}(\kappa),j_{X}(W_{X})\})$ (see Definition 5.14),
@@
(2.1) ${\sf{b}}$ b-catches $S_{\pi_{X}\circ\sigma^{\sf a}\,"\bar{\nu}}$ relative to $(P,j(\chi),\lambda,\{j(\kappa),j(W_{X})\})$,
@@
(2.3) $\pi_{X}\circ\sigma^{\sf a}\,"\bar{\nu}\subseteq j(W)$, and
@@
Assume there is a proper class of inaccessible limits of Woodin cardinals. Then $\mathsf{Generically}\ \allowbreak\mathsf{Correct}\ \allowbreak\mathsf{Sealing}$ implies $\mathsf{Sealing}$.
@@
such that for every $A\in\Gamma^{\infty}_{h_{1}}$, $j(A)=A_{h_{2}}$. The derived model theorem (see Theorem 2.2) implies that assuming $\mathsf{Sealing}$ and a class of inaccessible limits of Woodin cardinals, Clause (2) implies Clause (1) in Definition 1.8.